Preface

EuroForth is an annual conference on the Forth programming language, stack machines, and related topics, and has been held since 1985. The 24th Euro-Forth finds us in Vienna for the second time. The three previous EuroForths were held in Santander, Spain (2005), in Cambridge, England (2006), and in Schloss Dagstuhl, Germany (2007). Information on earlier conferences can be found at the EuroForth home page (http://dec.bournemouth.ac.uk/

forth/euro/index.html).

Since 1994, EuroForth has a refereed and a non-refereed track.

For the refereed track, one paper was submitted, and one was accepted (100% acceptance rate). For more meaningful statistics, I include the numbers from 2006 and 2007: eight submissions, four accepts, 50% acceptance rate. The paper was sent to three program committee members for review, and they produced three reviews. This year, one of the program committee members has submitted a paper; of course this member was not involved in the review process of the paper in any way. I thank the authors for their paper, and the reviewers for their reviews.

Several papers were submitted to the non-refereed track in time to be included in the printed proceedings. In addition, the printed proceedings include the slides and an abstract for talks that will be presented at the conference without being accompanied by a paper. These online proceedings also contain late papers and late presentations that were too late to be included in the printed proceedings. One presentation (GLforth by Gerald Wodni) was in demo format and is therefore not included in these proceedings. Workshops and social events complement the program.

We are grateful to Ewa Vesely for organizational support for this year’s EuroForth.

Anton Ertl

Program committee

Sergey N. Baranov, Motorola ZAO, Russia (secondary chair)

M. Anton Ertl, TU Wien (chair)

David Gregg, University of Dublin, Trinity College

Ulrich Hoffmann, FH Wedel University of Applied Sciences

Jaanus Pöial, Estonian Information Technology College, Tallinn Bradford Rodriguez, T-Recursive Technology Bill Stoddart, University of Teesside Reuben Thomas, Adsensus Ltd.

Contents

Programme . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .5

Refereed papers

Angel Robert Lynas, Bill Stoddart: Using Forth in a Concept-Oriented Computer Language Course . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

Non-refereed papers

Jaanus Pöial: Java Framework for Static Analysis of Forth Programs … 20 Stephen Pelc: Updating the Forth Virtual Machine . . . . . . . . . . . . . . . . . . . . . 24 Bernd Paysan: A Debugger for the b16 CPU . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 M. Anton Ertl: Cleaning up After Yourself . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

Abstracts

Ulrich Hoffmann: Forth System Hooks—Metaobject Protocol in Forth Style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39

Presentations

Yunhe Shi, Kevin Casey, M. Anton Ertl, David Gregg: Virtual Machine Showdown: Stack versus Registers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .40 Andrew Haley: An Infix Syntax for Forth (and its compiler) . . . . . . . . . . . . 48

Late papers

Bill Stoddart, Angel Robert Lynas: Filters for Unicode Markup, a Work in Progress Report . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

Late presentations

Bernd Paysan: Porting MINOS to VFX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 Federico de Ceballos: A high level interface to SQLite . . . . . . . . . . . . . . . . . . 66

Programme

Thursday 1:00pm Lunch at Artner
2:30pm Forth200x meeting
7:00pm Dinner at Wiazhaus
Friday 9:00am Forth200x meeting
1:00pm Lunch at Artner
2:30pm Coffee at meeting room
2:45pm EuroForth opening, organizational matters

3:00pm Angel Robert Lynas, Bill Stoddart: Using Forth in a

Concept-Oriented Computer Language Course

3:30pm Jaanus Pöial: Java Framework for Static Analysis of

Forth Programs

4:00pm Andrew Haley: An Infix Syntax for Forth (and its com-

piler)

4:30pm Coffee break
5:00pm Bernd Paysan: A Debugger for the b16 CPU
5:30pm M. Anton Ertl: Cleaning up After Yourself

6:00pm Ulrich Hoffmann: Forth System Hooks—Metaobject

Protocol in Forth Style

7:00pm Meet at hotel for walk to dinner
7:30pm Conference Dinner at Zwölfapostelkeller
Saturday 9:00am Stephen Pelc: Updating the Forth Virtual Machine

9:30am Yunhe Shi, Kevin Casey, M. Anton Ertl, David Gregg:

Virtual Machine Showdown: Stack versus Registers

10:00am Bernd Paysan: Porting MINOS to VFX
10:30am Coffee break

11:00am Bill Stoddart, Angel Robert Lynas: Filters for Unicode

Markup, a Work in Progress Report

11:30am Gerald Wodni: GLforth – an Ego Shooter
11:55am Impromptu talks/discussion
12:50pm Walk to restaurant
1:00pm Lunch at Wieden Bräu

2:30pm Excursion to Leopoldsberg, Kahlenberg, Cobenzl, fol-

lowed by Dinner at Weingut am Reisenberg

Sunday 9:00am EuroForth workshops and impromptu talks
1:00pm Lunch at Wieden Bräu
2:30pm 4th day: Walk: Alte Donau, Donauinsel and Prater